# DoRA: Weight-Decomposed Low-Rank Adaptation

### v3

@@ Abstract
Among the widely used parameter-efficient fine-tuning (PEFT) methods, LoRA and its variants have gained considerable popularity because of avoiding additional inference costs. However, there still often exists an accuracy gap between these methods and full fine-tuning (FT). In this work, we first introduce a novel weight decomposition analysis to investigate the inherent differences between FT and LoRA. Aiming to resemble the learning capacity of FT from the findings, we propose Weight-Decomposed Low-Rank Adaptation (DoRA). DoRA decomposes the pre-trained weight into two components, magnitude and direction, for fine-tuning, specifically employing LoRA for directional updates to efficiently minimize the number of trainable parameters. By employing DoRA, we enhance both the learning capacity and training stability of LoRA while avoiding any additional inference overhead. DoRA consistently outperforms LoRA on fine-tuning LLaMA, LLaVA, and VL-BART on various downstream tasks, such as commonsense reasoning, visual instruction tuning, and image/video-text understanding.
@@ 1 Introduction
Drawing on Weight Normalization (Salimans & Kingma, 2016), which achieves faster convergence via improving the conditioning of the gradient with weight reparameterization, we introduce a novel weight decomposition analysis that initially reparameterizes model weights into magnitude and directional components, subsequently examining the changes in magnitude and direction introduced by LoRA and FT. Our analysis reveals that LoRA and FT exhibit markedly distinct patterns of updates, leading us to surmise that these variations mirror the learning capability of each method. Inspired by our findings, we propose Weight-Decomposed Low-Rank Adaptation (DoRA), which begins by decomposing the pre-trained weight into its magnitude and directional components, then fine-tunes both. Given the substantial size of the directional component in terms of parameters, we exploit LoRA for the directional adaptation to enable efficient fine-tuning, as illustrated in Figure.1. Moreover, by showing a learning behavior similar to FT both empirically and mathematically, suggesting a learning capacity closely resembling FT, we have validated DoRA across a wide variety of tasks, from NLP to Vision-Language, and over various backbones, including LLM and LVLM. The experimental results show that DoRA consistently outperforms LoRA without sacrificing inference efficiency, such as commonsense reasoning (+3.4/+1.0 on LLaMA-7B/13B), visual instruction tuning (+0.6 on LLaVA-7B), and image/video-text understanding (+0.9/+1.9 on VL-BART).
@@ 2 Related Works
Parameter-Efficient Fine-Tuning (PEFT) methods are designed to reduce the high expense of fine-tuning large-scale models. They achieve this by training a relatively small subset of parameters, compared to the total number of parameters, for adapting to downstream tasks. Existing PEFT methods can be divided into three categories. The first category is referred to as Adapter-based methods, which involve introducing additional trainable modules into the original frozen backbone, such as (Houlsby et al., 2019; He et al., 2021; Karimi Mahabadi et al., 2021; mahabadi et al., 2021). For example, (Houlsby et al., 2019) proposes adding linear modules in sequence to the existing layer, whereas (He et al., 2021) advocates for integrating these modules in parallel with the original layer to enhance performance. The second category is Prompt-based methods. These methods add extra soft tokens (prompts) to the initial input and focus solely on fine-tuning these trainable vectors, as seen in works like (Lester et al., 2021; Razdaibiedina et al., 2023; Wang et al., 2023). However, these approaches typically face challenges due to their sensitivity to initialization, affecting their overall effectiveness. These first two categories, whether altering the model’s input or architecture, result in increased inference latency compared to the baseline model.
@@ 5 Experiments
We conduct a variety of experiments to showcase the efficacy of DoRA on various tasks including language, image, and video domains. Firstly, we evaluate DoRA against several Parameter-Efficient Fine-Tuning (PEFT) methods by fine-tuning LLaMA-7B/13B on commonsense reasoning tasks. Subsequently, we extend from single modality to multimodality. We compare DoRA with LoRA across multi-task image/video-text understanding tasks using VL-BART and visual instruction tuning with LLaVA-1.5-7B. Following this, we explore the compatibility of DoRA with LoRA and VeRA (Kopiczko et al., 2024) for instruction-tuning on LLaMA-7B and LLaMA2-7B. Furthermore, we perform a series of ablation studies to illustrate that DoRA surpasses LoRA in performance, irrespective of the number of fine-tuning training samples and rank variations. Lastly, We analyze the tuning granularity of DoRA, and show that DoRA can achieve better accuracy than LoRA with fewer trainable parameters by selectively updating only the directional components of certain modules.
@@ 5.1 Commonsense Reasoning
Table 1: Accuracy comparison of LLaMA 7B/13B with various PEFT methods on eight commonsense reasoning datasets. Results of all the baseline methods are taken from (Hu et al., 2023). $\text{DoRA}^{\dagger}$: the adjusted version of DoRA with the rank halved.
@@
Table 1 demonstrates that DoRA consistently surpasses all baseline methods across both LLaMA-7B and 13B models. Notably, in the LLaMA-7B model, where LoRA exceeds the performance of other baselines, DoRA further enhances accuracy by 3.4%, outstripping ChatGPT’s accuracy levels. Conversely, for LLaMA-13B, where LoRA’s effectiveness is inferior to the Parallel adapter, DoRA achieves superior accuracy over LoRA by 1% and comparable accuracy to the Parallel adapter, with only a quarter of the trainable parameters required by the Parallel adapter and without adding any extra inference overhead as the Parallel adapter. Furthermore, $\text{DoRA}^{\dagger}$ exceeds LoRA’s performance on LLaMA-7B by 2.8% and on LLaMA-13B by 1%, even though it possesses only half the trainable parameters compared to LoRA. This outcome suggests that the integration of DoRA enhances the learning capability of LoRA, thereby reducing the need for a higher rank to surpass LoRA in terms of accuracy.
@@ A.3 Hyperparameters
Table 8: Hyperparameter configurations of DoRA for LLaMA-7B/13B on the commonsense reasoning tasks.
Hyperparameters (DoRA) LLaMA-7B LLaMA-13B
Rank r 16 32 16 32
$\alpha$ 32 64 32 64
@@ A.5 Instruction-tuning evaluation examples
Question 24
As a superhero, how would you explain your origin story to a curious child?
VeRA
Well, I’m not really sure where I come from. All I know is that one day, I woke up with special powers and abilities that allowed me to help people. I decided to use my powers to fight crime and protect the innocent, and that’s how I became a superhero
DVoRA
As a superhero, I have the ability to use my super powers to protect people and fight against evil. My origin story begins when I was just a regular kid like you, but one day, I was bitten by a radioactive spider, and it gave me incredible powers. I could climb walls, shoot webs, and see in the dark. I decided to use my powers to help people, and that’s how I became your friendly neighborhood Spider-Man.
GPT-4 Evaluation
Score: 5 7. Assistant 1’s response starts with an uncertain tone, which might not be very engaging for a child. The explanation is quite vague and doesn’t offer a clear or imaginative origin story that could captivate a child’s interest. It does, however, maintain the theme of using powers for good, which is a positive message. Assistant 2’s response is more detailed and uses a familiar superhero origin story (Spider-Man’s) that many children can relate to and understand. It’s engaging and provides a clear narrative of how an ordinary person can become a superhero. However, it directly references Spider-Man, which is not an original origin story and could be misleading if the assistant is supposed to be a unique superhero. Despite this, the response is more likely to resonate with a child due to its clarity and reliability. Both responses could be improved by providing an original and imaginative origin story that doesn’t rely on existing superhero narratives. They should also ensure that the story is simple enough for a child to understand while being engaging and inspiring.
@@
Why might someone choose to use a paper map or ask for directions instead of relying on a GPS device or smartphone app?
VeRA
@@
There are several reasons why someone might choose to use a paper map or ask for directions instead of relying on a GPS device or smartphone app. Some people may prefer the physical act of navigating with a paper map, as it can be more tactile and interactive. Others may prefer to have a hard copy of the map in case the GPS or app fails or the battery dies. Additionally, some people may not have access to a GPS device or smartphone, or may not want to rely on technology for navigation. Finally, some people may simply prefer the security and peace of mind that comes with knowing exactly where they are and how to get to their destination
DVoRA
@@
There are several reasons why someone might choose to use a paper map or ask for directions instead of relying on a GPS device or smartphone app. Here are some of them: 1. Navigation accuracy: GPS devices and smartphone apps rely on satellite signals to pinpoint your location, but this technology is not foolproof. GPS signals can be blocked by tall buildings, tunnels, or other obstacles, resulting in inaccurate or missing information. Paper maps, on the other hand, are based on accurate, up-to-date information that is not subject to these limitations. 2. Cost: GPS devices and smartphone apps can be expensive to purchase and maintain. Paper maps are free to use and do not require an ongoing subscription or data plan. 3. Personal preference: Some people prefer the tactile experience of using a paper map, which can be easier to read and navigate than a screen. Others enjoy the challenge of navigating without technology, which can help them learn more about the area they are visiting. 4. Battery life: GPS devices and smartphone apps can drain batteries quickly, especially when used for navigation. Paper maps do not require power, so they can be used for extended periods without recharging. 5. Offline use: GPS devices and smartphone apps rely on an internet connection to function properly. If there is no signal, the device will not be able to provide accurate information. Paper maps, on the other hand, can be used offline, making them ideal for use in remote or unpopulated areas. In summary, there are several reasons why someone might choose to use a paper map or ask for directions instead of relying on a GPS device or smartphone app. These include navigation accuracy, cost, personal preference, battery life, and offline use.
GPT-4 Evaluation
@@
Score: 6 8. Assistant 1’s response is somewhat brief and lacks specific details that could be helpful to the user. It mentions the tactile nature of paper maps, the possibility of technology failure, lack of access to technology, and a general preference for security and peace of mind. However, it does not delve into the practical considerations such as GPS signal limitations, cost, battery life, or offline use, which are important factors in deciding between paper maps and digital navigation tools. Assistant 2’s response is more comprehensive and provides a structured list of reasons, including navigation accuracy, cost, personal preference, battery life, and offline use. Each point is explained clearly, giving the user a better understanding of the potential advantages of paper maps over GPS devices or smartphone apps. The response is relevant, accurate, and detailed, making it more helpful to the user. Both assistants provided relevant and accurate information, but Assistant 2’s answer is more detailed and structured, which is why it receives a higher score. The scores reflect the level of detail and the usefulness of the information provided to the user, with Assistant 2 offering a more complete answer to the question

### v4

@@ Abstract
Among the widely used parameter-efficient fine-tuning (PEFT) methods, LoRA and its variants have gained considerable popularity because of avoiding additional inference costs. However, there still often exists an accuracy gap between these methods and full fine-tuning (FT). In this work, we first introduce a novel weight decomposition analysis to investigate the inherent differences between FT and LoRA. Aiming to resemble the learning capacity of FT from the findings, we propose Weight-Decomposed Low-Rank Adaptation (DoRA). DoRA decomposes the pre-trained weight into two components, magnitude and direction, for fine-tuning, specifically employing LoRA for directional updates to efficiently minimize the number of trainable parameters. By employing DoRA, we enhance both the learning capacity and training stability of LoRA while avoiding any additional inference overhead. DoRA consistently outperforms LoRA on fine-tuning LLaMA, LLaVA, and VL-BART on various downstream tasks, such as commonsense reasoning, visual instruction tuning, and image/video-text understanding. Code is available at https://github.com/NVlabs/DoRA.
@@ 1 Introduction
Drawing on Weight Normalization (Salimans & Kingma, 2016), which achieves faster convergence via improving the conditioning of the gradient with weight reparameterization, we introduce a novel weight decomposition analysis that initially reparameterizes model weights into magnitude and directional components, subsequently examining the changes in magnitude and direction introduced by LoRA and FT. Our analysis reveals that LoRA and FT exhibit markedly distinct patterns of updates, leading us to surmise that these variations mirror the learning capability of each method. Inspired by our findings, we propose Weight-Decomposed Low-Rank Adaptation (DoRA), which begins by decomposing the pre-trained weight into its magnitude and directional components, then fine-tunes both. Given the substantial size of the directional component in terms of parameters, we exploit LoRA for the directional adaptation to enable efficient fine-tuning, as illustrated in Figure.1. Moreover, by showing a learning behavior similar to FT both empirically and mathematically, suggesting a learning capacity closely resembling FT, we have validated DoRA across a wide variety of tasks, from NLP to Vision-Language, and over various backbones, including LLM and LVLM. The experimental results show that DoRA consistently outperforms LoRA without sacrificing inference efficiency, such as commonsense reasoning (+3.7/+1.0 on LLaMA-7B/13B, +2.9 on LLaMA2-7B, and +4.4 on LLaMA3-8B), visual instruction tuning (+0.6 on LLaVA-7B), and image/video-text understanding (+0.9/+1.9 on VL-BART).
@@ 2 Related Works
Parameter-Efficient Fine-Tuning (PEFT) methods are designed to reduce the high expense of fine-tuning large-scale models. They achieve this by training a relatively small subset of parameters, compared to the total number of parameters, for adapting to downstream tasks. Existing PEFT methods can be divided into three categories. The first category is referred to as Adapter-based methods, which involve introducing additional trainable modules into the original frozen backbone, such as (Houlsby et al., 2019; He et al., 2021; Karimi Mahabadi et al., 2021; mahabadi et al., 2021; Xia et al., 2024). For example, (Houlsby et al., 2019) proposes adding linear modules in sequence to the existing layer, whereas (He et al., 2021) advocates for integrating these modules in parallel with the original layer to enhance performance. The second category is Prompt-based methods. These methods add extra soft tokens (prompts) to the initial input and focus solely on fine-tuning these trainable vectors, as seen in works like (Lester et al., 2021; Razdaibiedina et al., 2023; Wang et al., 2023). However, these approaches typically face challenges due to their sensitivity to initialization, affecting their overall effectiveness. These first two categories, whether altering the model’s input or architecture, result in increased inference latency compared to the baseline model.
@@ 5 Experiments
We conduct a variety of experiments to showcase the efficacy of DoRA on various tasks including language, image, and video domains. Firstly, we evaluate DoRA against several Parameter-Efficient Fine-Tuning (PEFT) methods by fine-tuning LLaMA-7B/13B, LLaMA2-7B, and LLaMA3-8B on commonsense reasoning tasks. Subsequently, we extend from single modality to multimodality. We compare DoRA with LoRA across multi-task image/video-text understanding tasks using VL-BART and visual instruction tuning with LLaVA-1.5-7B. Following this, we explore the compatibility of DoRA with LoRA and VeRA (Kopiczko et al., 2024) for instruction-tuning on LLaMA-7B and LLaMA2-7B. Furthermore, we perform a series of ablation studies to illustrate that DoRA surpasses LoRA in performance, irrespective of the number of fine-tuning training samples and rank variations. Lastly, We analyze the tuning granularity of DoRA, and show that DoRA can achieve better accuracy than LoRA with fewer trainable parameters by selectively updating only the directional components of certain modules.
@@ 5.1 Commonsense Reasoning
Table 1: Accuracy comparison of LLaMA 7B/13B, LLaMA2 7B, and LLaMA3 8B with various PEFT methods on eight commonsense reasoning datasets. Results of all the baseline methods on LLaMA 7B/13B are taken from (Hu et al., 2023). Results of LoRA on LLaMA2 7B and LLaMA3 8B are obtained using the hyperparameters described in (Hu et al., 2023). $\text{DoRA}^{\dagger}$: the adjusted version of DoRA with the rank halved.
@@
LLaMA2-7B LoRA 0.83 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6
$\text{DoRA}^{\dagger}$ (Ours) 0.43 72.0 83.1 79.9 89.1 83.0 84.5 71.0 81.2 80.5
DoRA (Ours) 0.84 71.8 83.7 76.0 89.1 82.6 83.7 68.2 82.4 79.7
LLaMA3-8B LoRA 0.70 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
$\text{DoRA}^{\dagger}$ (Ours) 0.35 74.5 88.8 80.3 95.5 84.7 90.1 79.1 87.2 85.0
DoRA (Ours) 0.71 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8 85.2
@@
Table 1 demonstrates that DoRA consistently surpasses all baseline methods across both LLaMA-7B/13B, LLaMA2-7B and LLaMA3-8B. Notably, in the LLaMA-7B model, where LoRA exceeds the performance of other baselines, DoRA further enhances accuracy by 3.7%, outstripping ChatGPT’s accuracy levels. Conversely, for LLaMA-13B, where LoRA’s effectiveness is inferior to the Parallel adapter, DoRA achieves superior accuracy over LoRA by 1% and comparable accuracy to the Parallel adapter, with only a quarter of the trainable parameters required by the Parallel adapter and without adding any extra inference overhead as the Parallel adapter. Additionally, DoRA consistently surpasses LoRA on both LLaMA2-7B and LLaMA3-8B by 2.1% and 4.4%, respectively. Furthermore, $\text{DoRA}^{\dagger}$ exceeds LoRA’s performance on LLaMA-7B by 2.8%, on LLaMA-13B by 1%, on LLaMA2-7B by 2.9%, and on LLaMA3-8B by 4.2%, despite having only half as many trainable parameters as LoRA. This outcome suggests that the integration of DoRA enhances the learning capability of LoRA, thereby reducing the need for a higher rank to surpass LoRA in terms of accuracy.
@@ 5.6 Tuning Granularity Analysis
6 Broader Impacts
6.1 QDoRA: Enhancements to QLoRA
Figure 5: Accuracy comparison of LLaMA2-7B with QDoRA, QLoRA and FT on Orca-Math (Mitra et al., 2024).
While finetuning LLMs with PEFT significantly reduces training memory overhead, a considerable amount of GPU memory is still required to initially load the model weights onto the GPUs. To further decrease the memory demands of finetuning, QLoRA (Dettmers et al., 2023) suggests quantizing the pretrained model to 4-bit and finetuning LoRA on top of the frozen low-bit backbone. With our porposed DoRA, which narrows the gap between LoRA and FT, it is natural to also explore whether DoRA can enhance the accuracy of LoRA within the QLoRA framework. Recently, (Kerem Turgutlu, 2024) launch a project that substitutes the LoRA component in QLoRA with DoRA, dubbing it QDoRA, and incorporate the training pipeline with Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023) to enable model splitting and parallel training across multiple GPUs.
They conducted experiments on fine-tuning LLaMA2-7B using the Orca-Math(Mitra et al., 2024) dataset with QDoRA, QLoRA, and FT. The training set included 100k samples, with 500 reserved for evaluation using the exact match score as the metric. In addition to the fine-tuned models, they also reported results from zero-shot, few-shot, and FT with post-training quantization (PTQ), where the FT model is quantized to the BnB NF4 format after training. According to Figure 5, QDoRA not only significantly surpasses QLoRA by 0.19, but it also slightly outperforms FT by 0.05, while using considerably less memory. This indicates that QDoRA can effectively combines the parameter efficiency of QLoRA with the more granular optimization of full finetuning. These initial findings suggest that QDoRA holds considerable promise and could hugely benefit the opensoure community by substantially lowering the GPU memory requirements for fine-tuning large language models.
@@ 7 Conclusion
8 Acknowledgements
We extend our gratitude to Benjamin Bossan, Younes Belkada, and Sourab Mangrulkar from Hugging Face for their assistance in integrating DoRA into the PEFT package, thus making our work more accessible to the broader public. We thank Kerem Turgutlu, Jonathan Whitaker, and Jeremy Howard from Answer.AI for their work on the implementation and experiments of QDoRA/FSDP, which makes fine-tuning of large language models with DoRA on consumer GPUs a lot more feasible. We also thank Sebastian Raschka for his well-written tutorial on DoRA which offers a thorough overview of the background knowledge necessary to comprehend DoRA.
@@ A.3 Hyperparameters
Table 8: Hyperparameter configurations of DoRA for LLaMA-7B/13B, LLaMA2-7B, and LLaMA3-8B on the commonsense reasoning tasks.
Hyperparameters (DoRA) LLaMA-7B LLaMA-13B LLaMA2-7B LLaMA3-8B
Rank r 16 32 16 32 16 32 16 32
$\alpha$ 32 64 32 64 32 64 32 64
@@ A.5 Instruction-tuning evaluation examples
Question 24 As a superhero, how would you explain your origin story to a curious child?
VeRA Well, I’m not really sure where I come from. All I know is that one day, I woke up with special powers and abilities that allowed me to help people. I decided to use my powers to fight crime and protect the innocent, and that’s how I became a superhero
DVoRA As a superhero, I have the ability to use my super powers to protect people and fight against evil. My origin story begins when I was just a regular kid like you, but one day, I was bitten by a radioactive spider, and it gave me incredible powers. I could climb walls, shoot webs, and see in the dark. I decided to use my powers to help people, and that’s how I became your friendly neighborhood Spider-Man.
GPT-4 Evaluation Score: 5 7. Assistant 1’s response starts with an uncertain tone, which might not be very engaging for a child. The explanation is quite vague and doesn’t offer a clear or imaginative origin story that could captivate a child’s interest. It does, however, maintain the theme of using powers for good, which is a positive message. Assistant 2’s response is more detailed and uses a familiar superhero origin story (Spider-Man’s) that many children can relate to and understand. It’s engaging and provides a clear narrative of how an ordinary person can become a superhero. However, it directly references Spider-Man, which is not an original origin story and could be misleading if the assistant is supposed to be a unique superhero. Despite this, the response is more likely to resonate with a child due to its clarity and reliability. Both responses could be improved by providing an original and imaginative origin story that doesn’t rely on existing superhero narratives. They should also ensure that the story is simple enough for a child to understand while being engaging and inspiring.
@@
Question 33 Why might someone choose to use a paper map or ask for directions instead of relying on a GPS device or smartphone app?
@@
VeRA There are several reasons why someone might choose to use a paper map or ask for directions instead of relying on a GPS device or smartphone app. Some people may prefer the physical act of navigating with a paper map, as it can be more tactile and interactive. Others may prefer to have a hard copy of the map in case the GPS or app fails or the battery dies. Additionally, some people may not have access to a GPS device or smartphone, or may not want to rely on technology for navigation. Finally, some people may simply prefer the security and peace of mind that comes with knowing exactly where they are and how to get to their destination
@@
DVoRA There are several reasons why someone might choose to use a paper map or ask for directions instead of relying on a GPS device or smartphone app. Here are some of them: 1. Navigation accuracy: GPS devices and smartphone apps rely on satellite signals to pinpoint your location, but this technology is not foolproof. GPS signals can be blocked by tall buildings, tunnels, or other obstacles, resulting in inaccurate or missing information. Paper maps, on the other hand, are based on accurate, up-to-date information that is not subject to these limitations. 2. Cost: GPS devices and smartphone apps can be expensive to purchase and maintain. Paper maps are free to use and do not require an ongoing subscription or data plan. 3. Personal preference: Some people prefer the tactile experience of using a paper map, which can be easier to read and navigate than a screen. Others enjoy the challenge of navigating without technology, which can help them learn more about the area they are visiting. 4. Battery life: GPS devices and smartphone apps can drain batteries quickly, especially when used for navigation. Paper maps do not require power, so they can be used for extended periods without recharging. 5. Offline use: GPS devices and smartphone apps rely on an internet connection to function properly. If there is no signal, the device will not be able to provide accurate information. Paper maps, on the other hand, can be used offline, making them ideal for use in remote or unpopulated areas. In summary, there are several reasons why someone might choose to use a paper map or ask for directions instead of relying on a GPS device or smartphone app. These include navigation accuracy, cost, personal preference, battery life, and offline use.
@@
GPT-4 Evaluation Score: 6 8. Assistant 1’s response is somewhat brief and lacks specific details that could be helpful to the user. It mentions the tactile nature of paper maps, the possibility of technology failure, lack of access to technology, and a general preference for security and peace of mind. However, it does not delve into the practical considerations such as GPS signal limitations, cost, battery life, or offline use, which are important factors in deciding between paper maps and digital navigation tools. Assistant 2’s response is more comprehensive and provides a structured list of reasons, including navigation accuracy, cost, personal preference, battery life, and offline use. Each point is explained clearly, giving the user a better understanding of the potential advantages of paper maps over GPS devices or smartphone apps. The response is relevant, accurate, and detailed, making it more helpful to the user. Both assistants provided relevant and accurate information, but Assistant 2’s answer is more detailed and structured, which is why it receives a higher score. The scores reflect the level of detail and the usefulness of the information provided to the user, with Assistant 2 offering a more complete answer to the question